Two point charges are separated by a distance of 40.0 cm. The numerical value of one charge is twice that of the other. Each charge exerts a force of magnitude 65.0 N on the other.

a.Find the magnitude of the charge with a smaller magnitude. (Express your answer to three significant figures.)

b.Find the magnitude of the charge with a greater magnitude. (Express your answer to three significant figures.)
Answer #1


We know, Force F = kq1q2 / r^2

65 N =9*10^9 * q * 2q / 0.2^2

Q=1.2* 10^-5 C

So, q1 = 1.2*10^-5 C (Ans)

And q2 = 2.4 * 10^-5 C (Ans)